• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ellie

    Log entry time 16:02:58 on October18,2010

    Entry number 335487

    keyword=Day Shift Summary

    The Day Shift was spent entirely in controlled access. The target warmed slightly this afternoon as Hall C fixed their moller, but it has been cooled back down to production mode. The LHRS is down (and has been all day) and needs to be brought back up before data taking can commence. The LHRS Q1 flow rate is currently being looked into by Heidi who will call shortly with an update.

    Good luck, swing shift!